# Benchmarking Resilience and Sensitivity of Polyurethane-Based Vision-Based Tactile Sensors

### v1

@@ Abstract
Vision-based tactile sensors (VBTSs) are a promising technology for robots, providing them with dense signals that can be translated into an understanding of normal and shear load, contact region, texture classification, and more. However, existing VBTS tactile surfaces make use of silicone gels, which provide high sensitivity but easily deteriorate from loading and surface wear. We propose that polyurethane rubber, used for high-load applications like shoe soles, rubber wheels, and industrial gaskets, may provide improved physical gel resilience, potentially at the cost of sensitivity. To compare the resilience and sensitivity of silicone and polyurethane VBTS gels, we propose a series of standard evaluation benchmarking protocols. Our resilience tests assess sensor durability across normal loading, shear loading, and abrasion. For sensitivity, we introduce model-free assessments of force and spatial sensitivity to directly measure the physical capabilities of each gel without effects introduced from data and model quality. Finally, we include a bottle cap loosening and tightening demonstration as an example where polyurethane gels provide an advantage over their silicone counterparts.
@@ I INTRODUCTION
The importance of tactile sensing in robotics is becoming increasingly recognized as robots transition from controlled laboratory environments to more structured and unpredictable scenarios like homes, factories, or elsewhere. Tactile sensing, the robotic equivalent of the human sense of touch, provides information about physical interactions with objects and the environment that allows robots to perform complex manipulation tasks, handle delicate objects, adapt to their surroundings, and recover from mistakes. Extensive research has been performed on the development and improvement of a range of tactile sensing technologies, including capacitive/piezoresistive/piezoelectric sensors, barometric sensors, and vision-based (or optical) tactile sensors [kappassov_tactile_2015, Cutkosky2016].
@@ I-A Vision Based Tactile Sensors
Among these technologies, vision-based tactile sensors (VBTSs) are increasingly promising due to the rich information they provide (e.g., Digit [lambeta_digit_2020], Gelsight [yuan_gelsight_2017]). These sensors operate by using an internal camera to observe the deformation of a soft elastomer that physically contacts the environment. The gel’s deformation is captured by an image transfer layer–often reflective and/or covered with a pattern of markers–which is cast onto a transparent base layer. This assembly is then adhered to an acrylic window in front of the internal camera [zhang_hardware_2022]. In existing work, both layers are typically made with silicone, a soft elastomer that provides high sensitivity during contact. VBTSs offer multiple advantages over more traditional tactile sensors, including high-resolution data, reduction of custom wires and circuitry, and the removal of electronics from the tactile interface itself [zhang_hardware_2022]. The visual data from the sensor can be processed to extract many types of information, such as contact region, normal and shear forces, texture classification, pose estimation, and slip detection [zhang_deltact_2022, baghaei_naeini_novel_2020, andrussow_minsight_2023, lin_dtact_2022, kakani_vision-based_2021, zhang_fingervision_2018].
Figure 1: Our polyurethane VBTS gel is capable of performing grasps without failure on rugged, heavy objects, including a 23.5 N piece of aluminum extrusion (left), a 40 N stool (center), and a 44.5 N weight (right).
@@ I-B Durability
However, one major limitation of VBTSs is the durability of their silicone gels. These gels have multiple known failure modes, including surface wear of the image transfer layer via abrasion/scratching/tearing and delamination of the silicone from the acrylic window [zhao_polytouch_2025, wang_gelsight_2021, potdar_high-speed_2024]. While their performance may be suitable for in-lab data collections and demonstrations, their deployment in the real world requires sensors that can endure repeated and unexpected use. Prior work has sought to improve the durability of gels by covering the sensor with a protective material, such as latex or tape [rayamane_design_2022, yamaguchi_combining_2016, wang_gelsight_2021, potdar_high-speed_2024]. However, these protective layers can cause artifacts in the image [wang_gelsight_2021, potdar_high-speed_2024] and often need replacing. Different works have also experimented with silicones of varying hardness [zhang_hardware_2022]. The low bond strength of silicone to acrylic, which leads to delamination failures, remains a problem across these approaches. As a result, most existing VBTS applications are currently limited to low-force manipulation tasks. In practice, these sensors may endure higher loads–possibly unexpectedly–and require an increased level of sensor resilience to various forms of loading and abrasion.
To address the need for a more resilient VBTS in home or industrial settings, we propose a new polyurethane-based gel. Polyurethane is used in various high-wear applications like shoe soles, skateboard wheels, and industrial gaskets, and may provide a significant lifespan improvement over its silicone counterparts. Polyurethane also bonds well to acrylic via superglue, which would reduce delamination failures. We predict that polyurethane will provide a more resilient alternative to silicone for VBTSs, although we expect that this will come at the cost of sensitivity.
I-C Benchmarking methods
@@
Limited work has formally assessed the durability of VBTS gels, and existing assessments focus only on a single, specific mode of failure [lambeta_digit_2020, rayamane_design_2022, zhao_polytouch_2025]. In reality, sensor gels will experience wear from a combination of different types of loading and rubbing. Additionally, while recent years have yielded many evaluations of VBTS performance [liu_tip_2025, higuera_sparsh_2024, schneider_tactile_2025], these evaluations are all model-based. As a result, their performance metrics depend on the specific model and dataset, which can obscure a direct understanding of the sensor’s intrinsic physical capabilities. While these model-based evaluations are useful for assessing performance in specific application contexts, we argue that a complementary, model-free evaluation can enable quicker, simpler baseline comparisons of sensitivity across different hardware platforms.
@@
To address these two concerns, we propose a collection of mechanical tests to evaluate sensitivity across a potential lifetime of wear, including compressive loading, shear loading, and abrasion. We also introduce model-free evaluations of sensitivity–specifically, force and spatial sensitivity–to allow for baseline comparisons across silicone and polyurethane VBTS gels. Our proposed spatial sensitivity evaluation uses frequency domain analysis to extract signal strength from a ridged surface with known period and amplitude. We take inspiration from grating orientation discrimination tasks, which have been used to assess tactile spatial sensitivity in humans [johnson_tactile_1981, van_boven_limit_1994, craig_grating_1999]. This task involves pressing a ridged surface onto a subject’s finger and examining whether they can correctly report its vertical or horizontal orientation. Recent work emulates this test, relying on the Structural Similarity Index Measure (SSIM) and a support vector machine (SVM) classifier to determine a tactile sensor’s ability to distinguish between vertical or horizontal ridges [liu_tip_2025]. However, this approach still relies on data collection and model training, and the use of SSIM (which reflects human visual perception) may present biases compared to the use of raw sensor data.
@@ I-D Overview
We present three novel contributions in this work:
@@
    The development of a polyurethane-based VBTS gel
@@
    VBTS mechanical resilience evaluations
  3. 3.
    VBTS model-free sensitivity evaluations
We begin by describing both our silicone and polyurethane gel fabrication techniques in Section II. Then, we present the parameters and procedures for our resilience and sensitivity evaluations in Section III. The results of these tests demonstrate that the polyurethane gel is more physically robust than the silicone gel, in Section IV, and outcomes are discussed in V. Finally, we include a repetitive bottle cap loosening and tightening demonstration in Section VI before concluding in Section VII.
@@ II GEL FABRICATION
In this section, we describe our fabrication techniques for producing the silicone and polyurethane gels used for comparison. We use the body of the DIGIT [lambeta_digit_2020] as a base, which contains a camera and three RGB LEDs. It uses a gel composed of a transparent silicone base layer (Smooth-On Solaris, Shore A 15) and an opaque image transfer layer (Smooth-On EcoFlex 00-10, Shore 00-10) colored with white pigment and airbrushed onto the base layer. This gel is glued onto an acrylic window with Smooth-On Sil-Poxy. The acrylic-gel unit can then be press-fit into the DIGIT case. Due to the DIGIT’s modularity and open-sourced design, we are able to develop and test custom acrylic-gel units with silicone and polyurethane. We choose to make our gels flat rather than rounded like the original DIGIT gels for ease of fabrication and better compatibility with our tests. We describe our manufacturing processes for each material below.
@@ II-1 Silicone Gels
We first create the silicone base layer with equal parts A and B (by weight) of Smooth-On Solaris. The mixture is degassed and then poured into a 4mm deep 3D printed mold to cure. To form the image transfer layer, we combine equal parts A and B of EcoFlex 00-10. We then mix in 10% by weight of Smooth-On Silicone Thinner, which reduces layer thickness for improved sensitivity, and 1% by weight of Smooth-On Cast Magic Silver Bullet powder and 2% by weight of Smooth-On Silc Pig white pigment, both of which contribute to lower opacity and higher reflectivity. To combine the base layer and image transfer layer, we place the cured base layer facing upwards on a flat acrylic plate, and the uncured image transfer layer mixture is poured over until it is covered. We found that this method produces an image transfer layer thicker than that of the original DIGIT, and thus less sensitive. To counter this, we use an air blow gun to gently displace some of the uncured material off the top of the base layer. While a more standardized procedure like spin-coating would provide better consistency, we find that our method is sufficient for small batches. After this step, we leave the silicone to cure for four hours. Once completely cured, we use a razor blade to clean up the edges of the gel before gluing it to a 6.35mm thick clear acrylic window with Smooth-On Sil-Poxy.
@@ II-2 Polyurethane Gels
To form the base layer of the polyurethane gel, we use equal parts by weight of Smooth-On Clear Flex 50 (Shore A 50) Parts A and B. We mix thoroughly and degas before pouring into a silicone negative mold created with a 3D printed positive and letting cure. The image transfer layer is created by mixing equal parts A and B of Smooth-On VytaFlex 40. To this mixture, we add 1% by weight of Smooth-On Cast Magic Silver Bullet powder and 2% by weight of Smooth-On SO-Strong white urethane colorant and mix thoroughly. To attach the base and image transfer layers, we first pour 0.25 g of the uncured image transfer layer mixture into the base of the emptied silicone mold. The specific volume of VytaFlex 40 is $971.8~\frac{\text{mm}^{3}}{\text{g}}$, yielding a layer height of approximately 0.62 mm (not considering mix-ins or mold fillets). We then use an air blow gun to gently spread the mixture along the base of the mold and remove large bubbles by popping with tweezers. Next, we place the base layer back into the mold and press it down onto the uncured image transfer layer, ensuring that no air is captured in pockets between the layers. Once placed evenly, the gel is left to cure. The cured gel is then attached to an acrylic window by spreading superglue across the surface of a window, pressing the gel on firmly, and wiping away excess glue.
Figure 3: Our model-free spatial sensitivity evaluation uses the sensor reading when pressed onto a ridged surface. The image is preprocessed via background subtraction and bandpass filtering before being cropped and run through a series of 1D FFTs. The resulting power spectral densities are averaged. The signal power is compared to a noise power, defined by the same frequency range from a flat, ridgeless surface, to obtain a signal-to-noise ratio (SNR). The background subtraction image is amplified for visualization purposes.
III SENSOR BENCHMARK METHODS
@@ III-A Mechanical Resilience Characterization
To characterize the mechanical resilience of our silicone and polyurethane gels, we put them through a series of mechanical tests that target different forms of loading or wear that may be encountered during use. As shown in Fig. 2, We include three categories of testing: compression, shear, and abrasion. We perform all tests with the sensor mounted to a Universal Robots UR-10 robot arm.
@@ III-A1 Cyclic Compression Loading
We first test the gel’s ability to withstand stretching or puncture from repeated compressive loading with the setup in Fig. 2A. Compressive loads may occur often during the lifetime of the sensor as it is used to grasp objects. We perform this test on a 4 mm spherical tip indenter with a 15 N compressive load and 1000 cycles. We selected a 15 N load through experimentation, increasing the load until failure was observed for either material. This loading is on the high end of typical human grasp forces [rajakumar_evidence_2022] but could easily be surpassed in more industrial settings [birglen_statistical_2018]. We test for 1000 cycles to approximate a lifetime of sensor use.
@@ III-A2 Cyclic Shear Loading
We also test the gel’s resilience under shear. We include both local shear and transverse shear, or shear across the whole face of the sensor. For local shear, we compress the sensor onto a 4 mm spherical tip indenter with a 10 N compressive load, and then apply a 5 N lateral load. This setup is shown in Fig. 2B. We choose a 10 N compressive load as it does not induce failure in either material alone within 1000 cycles, as revealed during the force selection process for the cyclic compression test. These loadings are similar to what humans might experience lifting a 500 g object with a small point of contact, assuming a static friction coefficient of $\sim$0.5. For transverse shear, we compress the sensor onto a flat acrylic plate with a 15 N compressive load and then apply a 15 N lateral load, as seen in Fig. 2C. We apply a 15 N compressive load to increase friction forces and prevent the gel from slipping on the plate. We include this test primarily to target delamination failures. The tests for local shear and shear across the whole gel are both run for 1000 cycles.

### v2

# Characterizing the Resilience and Sensitivity of Polyurethane Vision-Based Tactile Sensors

@@ Abstract
Vision-based tactile sensors (VBTSs) are a promising technology for robots, providing them with dense signals that can be translated into a multi-faceted understanding of contact. However, existing VBTS tactile surfaces make use of silicone gels, which provide high sensitivity but easily deteriorate from loading and surface wear. We propose that polyurethane rubber, a typically harder material used for high-load applications like shoe soles, rubber wheels, and industrial gaskets, may provide improved physical gel resilience, potentially at the cost of sensitivity. To compare the resilience and sensitivity of two polyurethane gel formulations against a common silicone baseline, we propose a series of repeatable characterization protocols. Our resilience tests assess sensor durability across normal loading, shear loading, and abrasion. For sensitivity, we introduce learning-free assessments of force and spatial sensitivity to directly measure the physical capabilities of each gel without effects introduced from data and model quality. We also include a bottle cap loosening and tightening demonstration to validate the results of our controlled tests with a real-world example. Our results show that polyurethane yields a more robust sensor. While it sacrifices sensitivity at low forces, the effective force range is largely increased, revealing the utility of polyurethane VBTSs over silicone versions in more rugged, high-load applications.
@@ I INTRODUCTION
The importance of tactile sensing in robotics is becoming increasingly recognized as robots transition from controlled laboratory environments to more structured and unpredictable scenarios like homes, factories, or elsewhere. Tactile sensing, the robotic equivalent of the human sense of touch, provides information about physical interactions with objects and the environment that allows robots to perform complex manipulation tasks, handle delicate objects, adapt to their surroundings, and recover from mistakes. Extensive research has been performed on the development and improvement of a range of tactile sensing technologies, including capacitive/piezoresistive/piezoelectric sensors, barometric sensors, and vision-based (or optical) tactile sensors [8, 4].
@@ I-A Vision Based Tactile Sensors
Among these technologies, vision-based tactile sensors (VBTSs) are increasingly promising due to the rich information they provide (e.g., Digit [9], Gelsight [19]). These sensors operate by using an internal camera to observe the deformation of a soft elastomer that physically contacts the environment. The gel’s deformation is captured by an image transfer layer–often reflective and/or covered with a pattern of markers–which is cast onto a transparent base layer. This assembly is then adhered to an acrylic window in front of the internal camera [21]. In existing work, both layers are typically made with silicone, a soft elastomer that provides high sensitivity during contact. VBTSs offer multiple advantages over more traditional tactile sensors, including high-resolution data, reduction of custom wires and circuitry, and the removal of electronics from the tactile interface itself [21]. The visual data from the sensor can be processed to extract many types of information, such as contact region, normal and shear forces, texture classification, pose estimation, and slip detection [20, 2, 1, 10, 7, 22].
Figure 1: The polyurethane VBTS gel (hardness 50A) is capable of performing grasps without failure on rugged, heavy objects, including a 23.5 N piece of aluminum extrusion (left), a 40 N stool (center), and a 44.5 N weight (right).
@@ I-B Durability
However, one major limitation of VBTSs is the durability of their silicone gels. These gels have multiple known failure modes, including surface wear of the image transfer layer via abrasion/scratching/tearing and delamination of the silicone from the acrylic window [23, 17, 12]. While their performance may be suitable for in-lab data collections and demonstrations, their deployment in the real world requires sensors that can endure repeated and unexpected use. Prior work has sought to improve the durability of gels by covering the sensor with a protective material, such as latex or tape [17, 12, 14, 18]. However, these protective layers can cause artifacts in the image [17, 12] and often need replacing. Different works have also experimented with silicones of varying hardness [21]. While this may improve resistance to tearing, delamination failures due to the low bond strength of silicone to acrylic remains a problem across all silicone versions. As a result, most existing VBTS applications are currently limited to low-force manipulation tasks. In practice, these sensors may endure higher loads–possibly unexpectedly–and require an increased level of sensor resilience to various forms of loading and abrasion.
Polyurethane, an elastomer used in various high-wear applications like shoe soles, skateboard wheels, and industrial gaskets, offers a promising alternative to silicone for improving VBTS durability. Subject to the requirement of transparency (for the base layer), readily available polyurethane rubbers have higher hardnesses than commonly used clear silicones, providing an intuitive advantage against mechanical wear. Polyurethane also bonds well to acrylic via superglue, which mitigates the delamination failure common to silicone versions. Despite these apparent advantages, the use of polyurethane for VBTS gels remains largely unexplored in current literature. As such, we predict that polyurethane will provide a more resilient alternative to typical silicones, but likely at the cost of force and spatial sensitivity.
I-C Sensor Characterization
@@
To characterize the resilience-sensitivity tradeoff across different gels, we develop a set of straightforward, repeatable evaluation protocols. We then use this framework to compare the performance of polyurethane sensors to a common silicone configuration as a baseline. Limited work has formally assessed the durability of VBTS gels, and existing assessments focus only on a single, specific mode of failure [9, 14, 23]. In reality, sensor gels will experience wear from a combination of different types of loading and rubbing. Additionally, while recent years have yielded many evaluations of VBTS performance [11, 5, 15], these evaluations are all learning-based. As a result, their performance metrics depend on the specific model and dataset, which can obscure a direct understanding of the sensor’s intrinsic physical capabilities. While these learning-based evaluations are useful for assessing performance in specific application contexts, we argue that a complementary, learning-free evaluation can enable quicker, simpler baseline comparisons of sensitivity across different hardware platforms.
@@
To address these two concerns, we propose a collection of mechanical tests to evaluate resilience across a potential lifetime of varying wear, including compressive loading, shear loading, and abrasion. We also introduce learning-free evaluations of sensitivity–specifically, force and spatial sensitivity–to allow for baseline comparisons across silicone and polyurethane VBTS gels. Our proposed spatial sensitivity evaluation uses frequency domain analysis to extract signal strength from a ridged surface with known period and amplitude. We take inspiration from grating orientation discrimination tasks, which have been used to assess tactile spatial sensitivity in humans [6, 16, 3]. This task involves pressing a ridged surface onto a subject’s finger and examining whether they can correctly report its vertical or horizontal orientation. Recent work emulates this test, relying on the Structural Similarity Index Measure (SSIM) and a support vector machine (SVM) classifier to determine a tactile sensor’s ability to distinguish between vertical or horizontal ridges [11]. However, this approach still relies on data collection and model training, and the use of SSIM (which reflects human visual perception) may present biases compared to the use of raw sensor data.
@@ I-D Overview
We present two novel contributions in this work:
@@
    The formal introduction and fabrication methodology of polyurethane VBTS gels for improved sensor resilience, and
@@
    A reproducible, learning-free sensor characterization framework for quantifying VBTS resilience (across compressive, shear, and abrasive wear) and force/spatial sensitivity.
We begin by describing both our baseline silicone and polyurethane gel fabrication techniques in Section II. Then, we present the parameters and procedures for our resilience and sensitivity characterizations in Section III. Results describing the tradeoff between resilience and sensitivity are detailed in Section IV, and outcomes are discussed in Section V. Finally, we include a bottle cap loosening and tightening demonstration in Section VI to validate resilience results in a real-world use case before concluding in Section VII.
@@ II GEL FABRICATION
In this section, we describe our fabrication techniques for producing the silicone and polyurethane gels used for comparison. We use the body of the DIGIT [9] as a base, which contains a camera and three RGB LEDs. It uses a gel composed of a transparent silicone base layer (Smooth-On Solaris, Shore A 15) and an opaque image transfer layer (Smooth-On EcoFlex 00-10, Shore 00-10) colored with white pigment and airbrushed onto the base layer. This gel is glued onto an acrylic window with Smooth-On Sil-Poxy. The acrylic-gel unit can then be press-fit into the DIGIT case. Due to the DIGIT’s modularity and open-sourced design, we are able to develop and test custom acrylic-gel units with silicone and polyurethane. We make our gels flat rather than rounded like the original DIGIT gels for ease of fabrication and better compatibility with our tests. We describe our manufacturing processes for each material below.
@@ II-1 Silicone Gels
For a silicone baseline, we replicate the material composition of the DIGIT gel due to its prevalence and similarity in hardness to other commonly used VBTSs like the GelSight family [21]. We first create the silicone base layer with equal parts A and B (by weight) of Smooth-On Solaris. The mixture is degassed and then poured into a 4mm deep 3D printed mold to cure. To form the image transfer layer, we combine equal parts A and B of EcoFlex 00-10. We then mix in 10% by weight of Smooth-On Silicone Thinner, which reduces layer thickness for improved sensitivity, and 1% by weight of Smooth-On Cast Magic Silver Bullet powder and 2% by weight of Smooth-On Silc Pig white pigment, both of which contribute to lower opacity and higher reflectivity. To combine the base layer and image transfer layer, we place the cured base layer facing upwards on a flat acrylic plate, and the uncured image transfer layer mixture is poured over until it is covered. We found that this method produces an image transfer layer thicker than that of the original DIGIT, and thus less sensitive. To counter this, we use an air blow gun to gently displace some of the uncured material off the top of the base layer. While a more standardized procedure like spin-coating would provide better consistency, we find that our method is sufficient for small batches. After this step, we leave the silicone to cure for four hours. Once completely cured, we use a razor blade to clean up the edges of the gel before gluing it to a 6.35mm thick clear acrylic window with Smooth-On Sil-Poxy.
@@ II-2 Polyurethane Gels
We fabricate two versions of a polyurethane gel using the same technique. The harder version uses Smooth-On Clear Flex 50 (Shore A 50) for the base layer and Smooth-On Vytaflex 40 (Shore A 40) for the reflective layer, while the softer version uses Smooth-On Clear Flex 30 (Shore A 30) and Smooth-On Vytaflex 20 (Shore A 20) in their place. We select these formulations since Clear Flex 30/50 are the softest readily available water-clear polyurethanes. To form the base layer of the polyurethane gel, we use the specified ratio of Clear Flex Parts A and B. We mix thoroughly and degas before pouring into a silicone negative mold created with a 3D printed positive and letting cure. The image transfer layer is created by mixing equal parts A and B of VytaFlex. To this mixture, we add 1% by weight of Smooth-On Cast Magic Silver Bullet powder and 2% by weight of Smooth-On SO-Strong white urethane colorant, mix thoroughly, and degas under vacuum. To attach the base and image transfer layers, we first pour 0.25 g of the uncured image transfer layer mixture into the base of the emptied silicone mold. The specific volume of VytaFlex 20/40 is $1000.7/971.8~\frac{\text{mm}^{3}}{\text{g}}$, yielding a layer height of approximately 0.64/0.62 mm (not considering mix-ins or mold fillets). We then use an air blow gun to gently spread the mixture along the base of the mold and remove large bubbles by popping with tweezers. Next, we place the base layer back into the mold and press it down onto the uncured image transfer layer, ensuring that no air is captured in pockets between the layers. Once placed evenly, the gel is left to cure. The cured gel is then attached to an acrylic window by spreading superglue across the surface of a window, pressing the gel on firmly, and wiping away excess glue.
Figure 3: (A) Our learning-free spatial sensitivity evaluation uses the sensor reading when pressed onto a ridged surface of varying period and amplitude. (B) The image is preprocessed via background subtraction and bandpass filtering before being cropped and run through a series of 1D FFTs. The resulting power spectral densities are averaged. The signal power is compared to a noise power, defined by the same frequency range from a flat, ridgeless surface, to obtain a signal-to-noise ratio (SNR). The background subtraction image is amplified for visualization purposes.
III SENSOR CHARACTERIZATION
We present a series of reproducible tests to characterize the resilience and sensitivity of each gel. The hardware necessary for each test is 3D-printable, and part files are available at github.com/[anonymized].
@@ III-A Mechanical Resilience Characterization
To characterize the mechanical resilience of our silicone and polyurethane gels, we put them through a series of mechanical tests representative of common wear modalities that may be encountered during use, including compression, shear, and abrasion (Fig. 2). For these tests, we select loading forces that 1) provide a good spread of data, and 2) are at or near those of human grasping and manipulation, a popular application for such sensors. In future work, reparameterization of these tests to represent loading conditions in different applications (e.g. medical or manufacturing) should be explored. We perform all tests with the sensor mounted to a Universal Robots UR-10 robot arm.
@@ III-A1 Cyclic Compression Loading
We first test the gel’s ability to withstand stretching or puncture from repeated compressive loading with the setup in Fig. 2A. Compressive loads may occur often during the lifetime of the sensor as it is used to grasp objects. We perform this test on a 4 mm spherical tip indenter with a 15 N compressive load and 1000 cycles. Testing with the narrow indenter, similar to grasping objects with sharper points or edges, provides a conservative estimate of resilience to compressive wear. The 15 N load is on the high end of typical human grasp forces [13] but was the minimum load where failure of any sensor observed. We test for 1000 cycles to reasonably approximate a lifetime of sensor use without requiring an extensive amount of time.
@@ III-A2 Cyclic Shear Loading
We also test the gel’s resilience under shear. We include both local shear and transverse shear, or shear across the whole face of the sensor. For local shear, we compress the sensor onto a 4 mm spherical tip indenter with a 10 N compressive load, and then apply a 5 N lateral load. This setup is shown in Fig. 2B. These loads are again representative of human grasp forces [13] and help to isolate the effect of local shear, as pilot testing revealed that all materials could withstand 10 N of cyclic compression alone. Local shear loads could occur when a grasped object with small contact points is subjected to unexpected external loads and may induce failures like tearing or puncture. For transverse shear, we compress the sensor onto a flat acrylic plate with a 15 N compressive load and then apply a 15 N lateral load, as seen in Fig. 2C. We apply a 15 N compressive load to increase friction forces and prevent the gel from slipping on the plate. This form of loading can cause delamination of sensor layers and is also important to investigate. The tests for local shear and shear across the whole gel are both run for 1000 cycles.